OpenStudy (anonymous):

Identify weather the given value of the variable makes the equation true. -5r-6=-26; r= 4

OpenStudy (zzr0ck3r):

does -5*4-6 equal 26?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

-20

OpenStudy (zzr0ck3r):

sorry I mean does -4*5-6 = -26?

OpenStudy (anonymous):

no I got -26

OpenStudy (anonymous):

oh sorry I thought u meant -27 LOL

OpenStudy (anonymous):

yeah it does

OpenStudy (zzr0ck3r):

so -5*4-6 = -26 so -5*r -6 = -26 when r is 4 so the given value of the variable makes the equation true.

OpenStudy (anonymous):

thanks

OpenStudy (zzr0ck3r):

no

OpenStudy (zzr0ck3r):

np*

OpenStudy (anonymous):

what about? 14x= -36 ; x = -3

OpenStudy (zzr0ck3r):

does 14*-3 = -36?

OpenStudy (anonymous):

no -42

OpenStudy (zzr0ck3r):

you have answered the question then

OpenStudy (anonymous):

so it doesnt ok thanks again